# Bio-Inspired Classification: Combining Information Theory and Spiking Neural Networks - Influence of the Learning Rules

### v1

@@ Abstract
Training of Spiking Neural Networks (SNN) is challenging due to their unique properties, including temporal dynamics, non-differentiability of spike events, and sparse event-driven activations. In this paper, we widely consider the influence of the type of chosen learning algorithm, including bioinspired learning rules on the accuracy of classification. We proposed a bioinspired classifier based on the combination of SNN and Lempel-Ziv complexity (LZC). This approach synergizes the strengths of SNNs in temporal precision and biological realism with LZC’s structural complexity analysis, facilitating efficient and interpretable classification of spatiotemporal neural data. It turned out that the classic backpropagation algorithm achieves excellent classification accuracy, but at extremely high computational cost, which makes it impractical for real-time applications. Biologically inspired learning algorithms such as tempotron and Spikprop provide increased computational efficiency while maintaining competitive classification performance, making them suitable for time-sensitive tasks. The results obtained indicate that the selection of the most appropriate learning algorithm depends on the trade-off between classification accuracy and computational cost as well as application constraints.
Index Terms:
spiking neural networks, learning algorithms, Artificial Intelligence, machine learning, networks dynamics, Shannon Information Theory, Lempel-Ziv Complexity, signal processing
I Introduction
@@
Backpropagation is a gradient-based learning rule commonly applied in networks training [1, 2]. Artificial Neural Networks (ANNs) and Spiking Neural Networks (SNNs) are both inspired by biological neural systems, however, they differ from the computation point of view. While ANNs dominate deep learning due to their simplicity, SNNs better mimic real neurons by using spike-based communication. However, training SNNs to match ANN accuracy, especially with backpropagation-based methods [1], requires significant hardware resources and is computationally expensive. As a consequence, SNNs training remains a challenge due to their complex dynamics and the non-differentiable nature of spike-based activity [3].
@@
Learning rules can be categorized into three main types: unsupervised learning (for example, Hebian learning [4], Symmetric Spike-Driven Synaptic Plasticity (SDSP) [5, 6]), and Spike Timing Dependent Plasticity (STDP) [8, 7]), supervised learning (like backpropagation (BP)[1, 2], Spike-Timing-Dependent Backpropagation (STBP) [10, 9], tempotron [13, 12, 11], SpikeProp [14, 15], Chronotron [16, 17], and Remote Supervised Method (ReSuMe) [18]), and hybrid learning (including Artificial-to-Spiking Neural Network conversion (ANN-SNN conversion) [19], and Reward-based STDP [20]).
@@
A commonly used training method in SNNs is SpikeProp, which is an arithmetic-based learning algorithm conceptually similar to the BP algorithm in ANNs [21]. However, arithmetic-based learning rules may be not a right choice for building biologically efficient networks. To achieve more biologically plausible learning, several alternative methods have been developed. For example, STDP that is inspired by biological synaptic plasticity, where neurons strengthen or weaken connections based on temporal contiguity. However, despite its efficiency in synaptic adaptation, STDP does not match the classification performance of BP-based ANN training [22, 23]. Moreover, spatio-temporal backpropagation (STBP) extends the traditional backpropagation algorithm to handle both spatial and temporal dependencies in spiking neural networks. This makes it possible to train deep SNNs using gradient-based optimization [24]. Other commonly used learning algorithms in SNN training are ReSuMe [18, 10, 25], and Chronotron [16, 17]. They are effective in adaptation to structured spike patterns, however, both are computationally demanding and require significant training time. An interesting SNN learning rule, namely tempotron was proposed by [13]. It is particularly effective for temporal coding, where information is embedded in the precise timing of spikes. In the study [26] training is conducted using the tempotron learning rule, while optimization is achieved through STDP. This combination significantly improves processing speed and reduces computational complexity compared to one can say "conventional" spiking neural network methods. In turn, [27, 28] proposed a neuron normalization technique and an explicitly iterative neuron model, which resulted in a significant increase in the SNNs’ learning rate. Another training strategy involves first training an ANN (typically using perceptrons) and then converting it into an SNN with an equivalent structure [19]. This conversion leverages the strengths of ANN training while allowing SNNs to benefit from energy-efficient spike-based processing. However, a major drawback is that this approach fails to capture temporal spike dynamics, as reliable frequency estimation requires a non-trivial time period. On the other hand, to optimize data efficiency, various active learning approaches have been introduced, including Bio-Inspired Active Learning (BAL), Bio-Inspired Active Learning based on Firing Rate (BAL-FR), and Bio-Inspired Active Learning utilizing Membrane Potential (BAL-M) [29]. These methods identify and select training samples by analyzing the discrepancies between empirical observations and generalization behavior, ensuring that previously unrecognized patterns are incorporated into the learning process. In turn, [9] introduces a Fractional-Order Spike-Timing-Dependent Gradient Descent (FO-STDGD) learning model, which applies fractional gradient descent principles to improve learning in SNNs.
We propose a novel hybrid approach that combines the temporal precision and biological plausibility of spiking neural networks with the Lempel-Ziv complexity (LZC) measure [41] to enhance classification performance. To our knowledge, this is the first time LZC has been integrated with SNNs. This method enables efficient and interpretable classification of spatiotemporal neural data by quantifying the structural complexity of spike patterns. It also improves robustness to noise and is particularly effective for data with variable temporal dynamics, such as Poisson-distributed signals. It turned out that the capability of bio-inspired neural networks in biosignal classification is highlighted, emphasizing their high precision, sensitivity, and reliability in distinguishing physiological patterns.
II Mathematical Formulation
II-A Notation
In Table I the basic notation is presented.
@@ II-A Notation
$\textbf{x}=[x_{1},x_{2},\cdots,x_{n}]\in\mathbb{R}^{n}$ synaptic weight vector
@@
$\tau_{m}\in\mathbb{R}^{+}$ membrane time constant
$t_{f}\in\mathbb{R}^{+}$ firing time of neuron
$S_{i}(t)\in\mathbb{Z}$ binary-valued spike train (0 or 1 at any given time)
$t_{i}^{k}\in\mathbb{R}^{+}$ spike times (real-valued, positive)
$\Theta(V_{n})\in\{0,1\}$ Heaviside function (1 if neuron fires, 0 otherwise)
$\eta\in\mathbb{R}^{+}$ general learning rate (positive real number) for weight update in supervised learning and gradient-based optimization
$\eta_{+},\eta_{-}\in\mathbb{R}^{+}$ STDP time constant (positive real)
$A\in\mathbb{R}$ SDSP learning rate (real-valuates), defines the strength of the symmetric weight update rule
$A_{-},A_{+}\in\mathbb{R}^{+}$ amplitude of STDP weight changes (positive real)
$t_{pre},t_{post}\in\mathbb{R}^{+}$ pre- and postsynaptic spike time (positive real)
$t_{spike}\in\mathbb{R}^{+}$ spike timing
$S_{pre}(t),S_{post}(t)\in\{0,1\}$ binary spike trains
$K(t)\in\mathbb{R}\longrightarrow\mathbb{R}$ kernel function, maps time to real value
$t_{i}^{target},t_{i}^{actual}\in\mathbb{R}^{+}$ targeted and ctual spike times
$E\in\mathbb{R}^{+}$ error function (positive real)
$\frac{\partial E}{\partial w_{i}}\in\mathbb{R}$ gradient of error
$\Phi(t_{pre},t_{post})\in\mathbb{R}^{+}$ loss function (positive real)
$S_{teach}(t)\in\{0,1\}$ binary teacher spike train
$w_{SNN},w_{ANN}\in\mathbb{R}$ synaptic weight in SNN/ANN (real-valued)
$\tau_{syn}\in\mathbb{R}^{+}$ synaptic time constant (positive real)
$r(t)\in\mathbb{R}$ reward signal (real-valued)
$\Delta w_{STDP}\in\mathbb{R}$ STDP-based weight change (real-valued)
$U(w_{i})\in\mathbb{R}\longrightarrow\mathbb{R}$ uncertainty function that maps weight to real value)
TABLE I: Basic notation used in the study.
II-B Unsupervised Learning Algorithms
Unsupervised learning algorithms learn patterns and structures from unlabeled data without explicit supervision or target output, i.e., provide local synaptic modifications with eplicit error signals.
II-B1 Hebbian Learning Rule
@@
II-B2 Spike Timing Dependent Plasticity
STDP is a biologically inspired rule that adjusts synaptic weights based on the relative timing of pre- and postsynaptic spikes [8, 7]:
@@ II-B2 Spike Timing Dependent Plasticity
It is a form of Hebbian learning (1). It is essential for modeling the temporal dynamics of biological neural systems. Unlike traditional learning methods, its event-driven nature enables unsupervised learning, allowing neural networks to adapt autonomously to temporal patterns in input data. This makes it valuable for processing spatio-temporal signals, such as sensory data. However, STDP requires careful parameter tuning, as factors like scaling and time constants significantly impact network behavior. In addition, its computational demands can be high, particularly in large-scale simulations [31].
II-B3 Symmetric Spike-Driven Synaptic Plasticity
SDSP is a variant of STDP, however, with symmetric weight updates [5, 6]
@@ II-B3 Symmetric Spike-Driven Synaptic Plasticity
It is a biologically plausible learning mechanism that mirrors neural dynamics by focusing on individual spike events. This event-driven approach enables rapid adaptability, making it valuable for real-time learning. However, its sensitivity to single spikes poses challenges in noisy environments, requiring advanced filtering techniques [32, 33].
II-C Supervised Learning Algorithms
Supervised learning maps input data to output labels based on labeled dataset. It relies on error-based updates.
II-C1 Backpropagation Algorithm
Backpropagation is a gradient-based learning rule commonly applied in networks training [1, 2].
@@ II-C1 Backpropagation Algorithm
II-C2 Spike-Timing-Dependent Backpropagation
Spike-Timing-Dependent Backpropagation extends the conventional gradient-based learning rule of backpropagation to spiking neural networks, incorporating spike-timing-dependent plasticity principles to optimize synaptic weights [10, 9]. The weight update rule is given by Equation (4). In STBP, the gradient computation follows the chain rule:
@@ II-C2 Spike-Timing-Dependent Backpropagation
II-C3 Tempotron Learning Rule
Tempotron learning rule adjust synaptics weight to classify temporal spike patterns [13, 12, 11]
@@ II-C3 Tempotron Learning Rule
The tempotron learning principle is a biologically inspired approach that refines synaptic weights based on the precise timing of incoming spikes rather than just their frequency. Unlike conventional neural models that prioritize weight adjustments or network structures, the tempotron underlines the importance of timing, showing that the "when" of a neural event can be as crucial as the "where" or "how often".
II-C4 SpikeProp Algorithm
SpikeProp applies gradient descent to minimize spike timing errors [14, 15]:

### v2

# arXiv:2506.06750v2[cs.AI] 01 Mar 2026

@@ Abstract
Training spiking neural networks (SNNs) remains challenging due to temporal dynamics, non-differentiability of spike events, and sparse event-driven activations. This paper studies how the choice of learning paradigm (unsupervised, supervised, and hybrid) affects classification performance and computational cost in temporal pattern recognition. Building on our earlier study [1], we use Lempel-Ziv complexity (LZC) as a compact, decision-relevant descriptor of spike-train temporal organization to quantify how different learning rules reshape class-conditional temporal structure. The pipeline combines a leaky integrate-and-fire (LIF) SNN with an LZC-based decision rule. We evaluate learning rules on synthetic sources with controlled temporal statistics (Bernoulli, two-state Markov, and Poisson spike processes) and on two-class subsets of MNIST and N-MNIST. Across datasets, gradient-based learning achieves the highest accuracy but at high computational cost, whereas bio-inspired rules (e.g., Tempotron and SpikeProp) offer favorable accuracy–efficiency trade-offs. These results highlight that selecting a learning rule should be guided by application constraints and the desired balance between separability and computational overhead.
keyword
spiking neural networks, learning rules, spike trains, temporal classification, Lempel-Ziv complexity, computational efficiency
1 Introduction
@@
Backpropagation is a gradient-based learning rule commonly applied in networks training [2]. Artificial Neural Networks (ANNs) and Spiking Neural Networks (SNNs) are both inspired by biological neural systems, however, they differ from the computation point of view. While ANNs dominate deep learning due to their simplicity, SNNs better mimic real neurons by using spike-based communication. However, training SNNs to match ANN accuracy, especially with backpropagation-based methods [2], requires significant hardware resources and is computationally expensive [3]. As a consequence, SNNs training remains a challenge due to their complex dynamics and the non-differentiable nature of spike-based activity [4, 5].
@@
Learning rules can be categorized into three main types: unsupervised learning (for example, Hebbian learning [6], Symmetric Spike-Driven Synaptic Plasticity (SDSP), and Spike Timing Dependent Plasticity (STDP) [8], supervised learning (like backpropagation (BP) [2], Spike-Timing-Dependent Backpropagation (STBP) [9], tempotron [10, 11], SpikeProp [12], its direct application to LIF [13] Chronotron [14], and Remote Supervised Method (ReSuMe [15]), and hybrid learning (including Artificial-to-Spiking Neural Network conversion (ANN-SNN conversion)) [16], and Reward-based STDP [8].
@@
A commonly used training method in SNNs is SpikeProp, which is an arithmetic-based learning algorithm conceptually similar to the BP algorithm in ANNs [17]. However, arithmetic-based learning rules may be not a right choice for building biologically efficient networks. To achieve more biologically plausible learning, several alternative methods have been developed. For example, STDP that is inspired by biological synaptic plasticity, where neurons strengthen or weaken connections based on temporal contiguity. However, despite its efficiency in synaptic adaptation, STDP does not match the classification performance of BP-based ANN training [2]. Moreover, spatio-temporal backpropagation (STBP) extends the traditional backpropagation algorithm to handle both spatial and temporal dependencies in spiking neural networks. This makes it possible to train deep SNNs using gradient-based optimization [18]. Other commonly used learning algorithms in SNN training are ReSuMe [15, 19], and Chronotron [14]. They are effective in adaptation to structured spike patterns, however, both are computationally demanding and require significant training time. An interesting SNN learning rule, namely tempotron was proposed by [10]. It is particularly effective for temporal coding, where information is embedded in the precise timing of spikes [20]. In the study [21] training is conducted using the tempotron learning rule, while optimization is achieved through STDP. This combination significantly improves processing speed and reduces computational complexity compared to one can say ”conventional” spiking neural network methods. In turn, [22] proposed a neuron normalization technique and an explicitly iterative neuron model, which resulted in a significant increase in the SNNs’ learning rate. Recently, the Spiking Gate (SG) ResNet architecture with an attention spike decoder (ASD) has been proposed to address gradient vanishing and improve spike decoding, achieving state-of-the-art performance on several benchmarks and enabling the first direct-training hybrid SNN-ANN detector [23]. Another training strategy involves first training an ANN (typically using perceptrons) and then converting it into an SNN with an equivalent structure [16]. This conversion leverages the strengths of ANN training while allowing SNNs to benefit from energy-efficient spike-based processing. However, a major drawback is that this approach fails to capture temporal spike dynamics, as reliable frequency estimation requires a non-trivial time period. On the other hand, to optimize data efficiency, various active learning approaches have been introduced, including Bio-Inspired Active Learning (BAL), Bio-Inspired Active Learning based on Firing Rate (BAL-FR), and Bio-Inspired Active Learning utilizing Membrane Potential (BAL-M) [24]. These methods identify and select training samples by analyzing the discrepancies between empirical observations and generalization behavior, ensuring that previously unrecognized patterns are incorporated into the learning process. In turn, [9] proposes the AC2AS framework, which couples activation consistency between ANN and SNN to enable fast and memory-efficient training of spiking neural networks.
We use a previously introduced LZC-SNN analysis framework [25] to study how learning rules reshape temporal spike-train organization. In this work, LZC is treated as a compact descriptor of temporal novelty and regularity, and the downstream classification performance is interpreted as a probe of separability induced by learning rather than as evidence of a universally strong stand-alone classifier. It turned out that the capability of bio-inspired neural networks in biosignal classification is highlighted, emphasizing their high precision, sensitivity, and reliability in distinguishing physiological patterns.
This study builds explicitly upon our earlier work [1], in which Lempel–Ziv complexity (LZC) was introduced as an analysis tool to investigate how different neuron models affect spiking neural network dynamics and downstream separability. In contrast to that work, the present manuscript shifts the scientific focus from neuron-model selection to the role of learning algorithms themselves.
Rather than re-introducing the coupling between LZC and spiking neural networks, we use LZC here as a compact temporal descriptor to quantify how unsupervised, supervised, and hybrid learning paradigms reshape the temporal organization of spike trains. This shift allows us to analyze learning rules not only in terms of classification accuracy, but also in terms of how they reorganize temporal variability, correlations, and computational efficiency under controlled input statistics.
While the underlying SNN architecture and part of the synthetic data generation are shared between the two studies, the scope and objectives differ substantially. Here, LZC is not used merely as a descriptive post hoc metric but as a decision-relevant representation to analyze the effect of unsupervised, supervised, and hybrid learning rules on class-conditional temporal complexity, robustness, and computational cost.
In this study, we address the problem of comparing learning rules in spiking neural networks from the perspective of temporal organization and computational efficiency. We propose a unified experimental protocol covering unsupervised, supervised, and hybrid learning paradigms under controlled temporal statistics, complemented by two-class MNIST and N-MNIST subsets. Within this framework, we demonstrate that learning rules with similar accuracy can produce substantially different class-conditional Lempel-Ziv complexity profiles, highlighting differences in temporal structure that are not reflected in accuracy metrics alone. By jointly analyzing accuracy and computational overhead, we identify operating regimes in which biologically inspired learning rules provide favorable efficiency-accuracy trade-offs. We also discuss the limitations of scalar complexity-based decisions, particularly under class imbalance and overlapping complexity distributions, and position Lempel-Ziv complexity as a decision-relevant descriptor rather than a universal classifier.
1.1 Relation to Prior Work
Table 1 summarizes the key differences between the present study and our earlier study [1]. Although both studies employ a similar SNN architecture and synthetic spike-based datasets, the scientific questions, evaluation criteria, and conclusions are fundamentally different. The earlier study focused on the influence of neuron models on SNN dynamics, whereas the present work isolates the role of learning algorithms and their impact on temporal information structure, robustness, and computational efficiency.
Table 1: Comparison between the present study and prior work [1].
Aspect Neuroinformatics (2026) [1] This work
Primary focus Neuron model dynamics Learning rule dynamics
Role of LZC Descriptive analysis metric Decision-relevant temporal descriptor
Learning algorithms Fixed learning setup Unsupervised, supervised, and hybrid learning
Evaluation criteria Accuracy and neuron behavior Accuracy, LZC profiles, and computational cost
Scientific question How neuron models affect SNNs How learning rules reshape temporal structure
To ensure transparency, we note that the network topology and part of the synthetic data generation follow [1]. However, the present study introduces a different experimental axis (learning rules), new evaluation dimensions (computational cost and complexity profiles under training), and additional datasets to assess generalization under controlled constraints.
Importantly, although both studies share a comparable LIF-based network topology and part of the synthetic data generation, the evaluated object is fundamentally different. The earlier work isolated the impact of neuron models under a fixed learning setup, whereas the present study isolates the impact of learning rules under a fixed neuron model. Consequently, the reported effects concern different mechanisms: here we focus on how learning dynamics reshape temporal correlations and variability in spike trains, as exposed by shifts in class-conditional LZC distributions and accuracy-efficiency trade-offs.
2 Background on Learning Rules in Spiking Neural Networks
Rather than treating learning rules solely as optimization procedures, we interpret them as mechanisms that actively shape the temporal structure of spike trains. Since Lempel-Ziv complexity directly quantifies the emergence of novel temporal patterns, different learning paradigms are expected to induce distinct complexity profiles even when classification accuracy is similar. In this work, learning algorithms are therefore grouped according to how they modulate temporal correlations, variability, and information flow in spiking activity, which in turn affects complexity-based pattern recognition.
In this section, we provide the background on the learning rules considered in this study and introduce the notation used throughout the paper. We first summarize the symbols and variables employed in the remainder of the manuscript (Section 2.1). We then group the learning algorithms for spiking neural networks into three main families: unsupervised, supervised, and hybrid rules. For each family, we briefly recall the underlying learning principle, present the corresponding mathematical formulation, and highlight its main advantages and limitations in the context of temporal pattern recognition. This structured overview prepares the ground for the description of the classification task and the proposed LZC-based SNN classifier in the subsequent sections.
2.1 Notation
In Table 2 the basic notation is presented.
@@ 2.1 Notation
$\mathbf{x}=[x_{1},x_{2},\dots,x_{n}]\in\mathbb{R}^{n}$ Synaptic weight vector
@@
$t_{f}\in\mathbb{R}^{+}$ Neuron firing time
$S_{i}(t)\in\{0,1\}$ Binary spike indicator at time $t$
$t_{i}^{k}\in\mathbb{R}^{+}$ Time of the $k$-th spike of neuron $i$
$H(V_{m}(t))\in\{0,1\}$ Heaviside step function
$\eta\in\mathbb{R}^{+}$ Learning rate
$\tau_{+},\tau_{-}\in\mathbb{R}^{+}$ STDP time constants
$A\in\mathbb{R}$ SDSP learning rate
$A_{+},A_{-}\in\mathbb{R}^{+}$ STDP amplitudes
$S_{\mathrm{pre}}(t),S_{\mathrm{post}}(t)\in\{0,1\}$ Pre- and postsynaptic binary spike trains
$t_{\mathrm{pre}},t_{\mathrm{post}}\in\mathbb{R}^{+}$ Pre- and postsynaptic spike times
$K(t):\mathbb{R}\rightarrow\mathbb{R}$ Synaptic kernel used in spike-timing rules (e.g., Tempotron)
$r(t)\in\mathbb{R}$ Reward signal in reward-modulated plasticity
$w_{\mathrm{SNN}},w_{\mathrm{ANN}}\in\mathbb{R}$ Synaptic weight in SNN / ANN (conversion setting)
$\tau_{\mathrm{syn}}\in\mathbb{R}^{+}$ Synaptic time constant
$U(w_{i}):\mathbb{R}\rightarrow\mathbb{R}$ Uncertainty function used in active learning updates
Table 2: Basic notation used in this study.
Unsupervised learning algorithms learn patterns and structures from unlabeled data without explicit supervision or target output, i.e., provide local synaptic modifications without explicit error signals. Unsupervised learning rules in spiking neural networks adjust synaptic weights based solely on the statistics and timing of pre- and postsynaptic activity, without access to explicit target labels. They provide local, biologically plausible synaptic updates and are particularly suitable for capturing regularities in continuous, unlabeled spike streams. In this work we focus on three widely used unsupervised mechanisms, Hebbian learning, spike-timing-dependent plasticity, and symmetric spike-driven synaptic plasticity, as they form a representative set of bio-inspired rules against which we later compare supervised and hybrid approaches. From an information-theoretic perspective, correlation-driven unsupervised rules tend to suppress temporal variability by reinforcing frequently co-occurring spike patterns. As a consequence, they often lead to reduced LZC values and narrower complexity distributions, which can limit discriminative power for inputs with overlapping temporal statistics.
2.2 Unsupervised Learning Algorithms
@@
2.2.1 Spike Timing Dependent Plasticity
Symmetric spike-driven plasticity refers to a class of local learning rules with symmetric weight updates driven by spike events, rather than a single canonical algorithm
@@ 2.2.1 Spike Timing Dependent Plasticity
It is a form of Hebbian learning (1). It is essential for modeling the temporal dynamics of biological neural systems. Unlike traditional learning methods, its event-driven nature enables unsupervised learning, allowing neural networks to adapt autonomously to temporal patterns in input data. This makes it valuable for processing spatio-temporal signals, such as sensory data. In addition, its computational demands can be high in large-scale simulations.
2.2.2 Symmetric Spike-Driven Spike-Based Plasticity
SDSP is a variant of STDP, however, with symmetric weight updates
@@ 2.2.2 Symmetric Spike-Driven Spike-Based Plasticity
It is a biologically plausible learning mechanism that mirrors neural dynamics by focusing on individual spike events. This event-driven approach enables rapid adaptability, making it valuable for real-time learning. However, its sensitivity to single spikes poses challenges in noisy environments, requiring advanced filtering techniques.
2.3 Supervised Learning Algorithms
Supervised learning maps input data to output labels based on labeled dataset. It relies on error-based updates. They adjust synaptic weights based on an explicit error or teacher signal that encodes the desired network output. In contrast to unsupervised rules, they are designed to optimize task-specific performance, such as classification accuracy, at the cost of reduced biological plausibility and typically higher computational complexity. Here we review representative supervised learning rules for SNNs, including gradient-based methods (backpropagation and STBP) and spike-timing-based algorithms (tempotron, SpikeProp, Chronotron, and ReSuMe), all of which are later evaluated within the same classification framework. Error-driven learning rules explicitly optimize class separability, which manifests not only in improved accuracy but also in increased divergence between class-conditional LZC distributions. This effect highlights that supervised learning influences not only decision boundaries but also the internal temporal organization of spike trains.
2.3.1 Backpropagation Algorithm
Backpropagation is a gradient-based learning rule commonly applied in networks training [2].
@@ 2.3.1 Backpropagation Algorithm
2.3.2 Spike-Timing-Dependent Backpropagation
Spike-Timing-Dependent Backpropagation extends the conventional gradient-based learning rule of backpropagation to spiking neural networks, incorporating spike-timing-dependent plasticity principles to optimize synaptic weights [9]. The weight update rule is given by Equation (4). In STBP, the gradient computation follows the chain rule:
@@ 2.3.2 Spike-Timing-Dependent Backpropagation
2.3.3 Tempotron Learning Rule
Tempotron learning rule adjust synaptics weight to classify temporal spike patterns [10, 29]
@@ 2.3.3 Tempotron Learning Rule
The tempotron learning principle is a biologically inspired approach that refines synaptic weights based on the precise timing of incoming spikes rather than just their frequency. Unlike conventional neural models that prioritize weight adjustments or network structures, the tempotron underlines the importance of timing, showing that the ”when” of a neural event can be as crucial as the ”where” or ”how often”.
2.3.4 SpikeProp Algorithm
SpikeProp applies gradient descent to minimize spike timing errors [12]: